OpenStudy (anonymous):

Given the function f(x) = 2(x + 10), find x if f(x) = 24. 68 7 17 2

OpenStudy (anonymous):

68

OpenStudy (anonymous):

how did yu do this

OpenStudy (anonymous):

2(24+10) = 48+20 68

OpenStudy (anonymous):

\(f(x)=24 \Rightarrow 2(x+10)=24\) Solve for \(x\).

OpenStudy (anonymous):

@desibaybee - That is incorrect. You found \(f(24)\). Instead, find \(24=f(x)=2(x+10)\), which means to solve \(24=2(x+10)\).

OpenStudy (anonymous):

2x+20=24

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

Ok LOL SO THE ANSWER IS 2 !!!!

OpenStudy (anonymous):

Whoopsss! sorry

OpenStudy (anonymous):

2(x+10)=24 2x +20=24 -20 -20 2x=4/2 x=2

OpenStudy (anonymous):

@freddie11 So you got that? :)